OpenStudy (anonymous):

Medals Given Solve 6 1/9 - 3 1/3. Show your work. But write it out.

OpenStudy (anonymous):

9006 approximately because 5x + 10 - sqroot 15 = 9005.6

OpenStudy (♪chibiterasu):

Wow pete you were waaaaay off o-o We just need to convert eh 3rd into a 8th, peaches @_@ So multiplication by 3 will help, yes it will :3 1/3 • 3/3 = 3/9 Now we have: 6 1/9 - 3 3/9 6 - 3 = 3 1 - 3 = -2 3 -2/9 Which simplifies to: -0.67 repeated. (66 repeats)

OpenStudy (♪chibiterasu):

@charlotte123 smack dis boi ;-;

OpenStudy (anonymous):

i disagree, you forgot to apply newton's method of root approximation, which would add an additional 5x^e, making it 9006.

OpenStudy (anonymous):

were just doing simple math

OpenStudy (charlotte123):

@♪Chibiterasu *Prepares hand....................*.....................*Runs* D:

OpenStudy (♪chibiterasu):

Pete over-thinks 5th grade math! It's not very effective ;-; Chibi used logic. It's super effective! :D pete gets shunned, Chibi wins the battle.

OpenStudy (charlotte123):

\(\huge\color{black}{\frak{For~One~Cannot~Simply~Defeat~The~Chi}}\)

OpenStudy (♪chibiterasu):

^^ Truth man ಠ_ಠ